7. (10 points) Find the general solution to the homogeneous system of DE: x = Ax where A = [-2 21
Answer #1

20 -2 tu) -2 -11 (lting 2 + 2x2 20 2 1² + 24 + 520 » N2 -2 + √4-20 2 V-16 2 = -1 + 2i - e-t/cosfat) & sin(2+)] 7 x= AX Where